### v1

# Breaking the Reviewer: Assessing the Vulnerability of Large Language Models in Automated Peer Review Under Textual Adversarial Attacks

Abstract
Peer review is essential for maintaining academic quality, but the increasing volume of submissions places a significant burden on reviewers. Large language models (LLMs) offer potential assistance in this process, yet their susceptibility to textual adversarial attacks raises reliability concerns. This paper investigates the robustness of LLMs used as automated reviewers in the presence of such attacks. We focus on three key questions: (1) The effectiveness of LLMs in generating reviews compared to human reviewers. (2) The impact of adversarial attacks on the reliability of LLM-generated reviews. (3) Challenges and potential mitigation strategies for LLM-based review. Our evaluation reveals significant vulnerabilities, as text manipulations can distort LLM assessments. We offer a comprehensive evaluation of LLM performance in automated peer reviewing and analyze its robustness against adversarial attacks. Our findings emphasize the importance of addressing adversarial risks to ensure AI strengthens, rather than compromises, the integrity of scholarly communication.
Breaking the Reviewer: Assessing the Vulnerability of Large Language Models in Automated Peer Review Under Textual Adversarial Attacks
Tzu-Ling Lin1, Wei-Chih Chen1, Teng-Fang Hsiao1, Hou-I Liu1, Ya-Hsin Yeh1, Yu Kai Chan1, Wen-Sheng Lien1, Po-Yen Kuo1, Philip S. Yu2, Hong-Han Shuai1thanks: Corresponding author. 1National Yang Ming Chiao Tung University 2University of Illinois Chicago
1 Introduction
Peer review is crucial for examining the quality and integrity of academic publications. However, recent studies Pan et al. (2018); Fortunato et al. (2018); Bornmann et al. (2021); Weissgerber et al. (2021); Severin and Chataway (2021); Shah (2022) highlighted a rapid increase in manuscript submissions, placing a significant burden on reviewers. Increasing specialization within fields further complicates the process, making it harder for reviewers to provide insightful comments on topics outside their expertise. Consequently, some reviews are less comprehensive or insightful than they might be with more time or specialized knowledge Shah (2022); Shah et al. (2018); Severin and Chataway (2021); Staudinger et al. (2024).
In response to these challenges, there is growing interest in leveraging automated systems to assist in the peer review process. Advances in natural language processing (NLP) and machine learning have opened promising avenues for developing tools that can augment human efforts in review generation. Early work by Nikiforovskaya et al. (2020) proposed an automated review generation process that combines traditional bibliometric analysis with BERT-based extractive summarization to identify key contributions in manuscripts. More recently, Zhou et al. (2024); Du et al. (2024) conducted a comprehensive evaluation of LLMs on automatic paper reviewing tasks, exploring the reliability of LLMs as potential reviewers.
Despite the benefits, the integration of LLMs into the peer review process raises concerns. One pressing issue is the vulnerability of LLM-generated reviews to textual adversarial attacks. Given that LLMs heavily rely on textual input, simple manipulations of the content could significantly distort their assessments. This vulnerability poses a security risk, especially if reviewers, intentionally or unintentionally, use mainstream LLMs to generate reviews without rigorous oversight.
In this paper, we investigate the robustness of LLMs as scientific paper reviewers. Specifically, we address three research questions:
(1) How effectively can LLMs generate reviews that match the quality of human-written ones?
(2) To what extent do textual adversarial attacks impact the reliability of LLM-generated reviews?
(3) What challenges arise when applying textual adversarial attacks to LLM-based reviewing systems, and how can they be mitigated?
To address these questions, we begin by evaluating the performance of leading LLMs, including closed-source models (e.g., GPT-4o, GPT-4o-mini), as well as open-source models (e.g., Llama-3.3-70B, Mistral-Small-3.1), on two core peer review tasks: review generation and score prediction. Our evaluation demonstrates the model’s capability to produce coherent and fluent reviews across multiple assessment aspects. We then apply common textual adversarial attack techniques to test the robustness of the LLM’s evaluations. Our findings reveal significant vulnerabilities, as the LLM’s assessments can be easily distorted through simple manipulations of the input text. Our contributions are threefold:
  • Aspect-based evaluation of LLM reviewer: We provide an in-depth analysis of an LLM’s performance in automatic paper reviewing tasks, comparing its outputs to those of human reviewers using aspect-based score.
  • Adversarial robustness analysis: We assess the impact of textual adversarial attacks on the reliability of LLM-generated reviews, highlighting specific vulnerabilities.
  • Limitations of existing defenses: We analyze current defense methods and show that they struggle to detect adversarial manipulations or reliably identify LLM-generated reviews for securing the fairness of review systems.
2 Related Work
2.1 Automated Paper Review-Related Tasks
The surge in academic publications has overwhelmed traditional peer review systems Weissgerber et al. (2021); Severin and Chataway (2021); Shah (2022), prompting the development of automated tools for tasks like readability assessment, plagiarism detection, and statistical error checking Kilicoglu et al. (2018); Riedel et al. (2020, 2022); Zhang (2010); Nuijten et al. (2016); Checco et al. (2021); Schulz et al. (2022). While helpful, these tools lack the ability to evaluate scientific content or methodological soundness.
Recent advances in NLP enable deeper automation. Nikiforovskaya et al. (2020) and Yuan et al. (2022) proposed automatic review generation pipelines using BERT and BART, respectively. Lu et al. (2024) introduced a framework for LLM-driven scientific research and evaluation. Du et al. (2024) investigated LLMs as (meta-)reviewers in NLP, while Yu et al. (2024a) proposed a framework combining review generation and self-critique.
Several works Liu and Shah (2023); Robertson (2023); Zhou et al. (2024); Liang et al. (2024) assess GPT-4’s review quality, revealing its potential and limitations. Other studies Wu et al. (2025, 2024); Yu et al. (2024b) address the detectability of LLM-generated reviews, emphasizing the difficulty of avoiding false positives and negatives, which complicates efforts to safeguard review integrity.
2.2 Vulnerabilities of LLMs to Adversarial Text Attacks
Adversarial text attacks manipulate inputs at different levels—characters Gao et al. (2018); Ebrahimi et al. (2018); Belinkov and Bisk (2017); Wei et al. (2024), words Jin et al. (2020); Li et al. (2020); Maheshwary et al. (2021), or sentences Qi et al. (2021a, b)—to mislead models without altering the core meaning. These attacks, whether white-box or black-box, expose vulnerabilities in LLMs with minimal and often imperceptible changes.
Recent surveys Yao et al. (2024); Kumar (2024) highlight growing concerns, including data poisoning, prompt injection, and biases such as position and verbosity bias Liu et al. (2024); Saito et al. (2023), or self-enhancement bias Zheng et al. (2023). In peer review, Robertson (2023) and Raina et al. (2024) show that subtle adversarial edits can mislead GPT-based reviewers, inflating scores or obscuring flaws, underscoring the risks in relying solely on LLMs for evaluation tasks.
3 LLM Vulnerability on Paper Reviews
Figure 1: The scenario illustration for LLM under adversarial attack, which author can be potential attacks introduce specific patterns (e.g., typos) into paper content. Thus, the feedback generated by LLM reviewers was misleading.
In response to the growing use of LLMs in peer review Zhou et al. (2024); Tyser et al. (2024); Yu et al. (2024b); Lu et al. (2024), we investigate their ability to automatically generate scientific reviews and their robustness against textual adversarial attacks—subtle edits that preserve meaning but may influence model outputs.
In Section 3.1, we simulate a realistic review workflow in which reviewers employ LLMs in a zero-shot manner to produce aspect-tagged reviews and scores with minimal manual effort. To evaluate the resilience of this process, we introduce Attack Focus Localization in Section 3.2, a method that identifies effective textual adversarial attack candidates in extremely long inputs via Longest Common Subsequence (LCS) matching. In Section 3.3, we integrate this localization algorithm with existing adversarial perturbations (e.g., typos or pattern-based edits at the character, word, or sentence level), which can potentially bias the LLM’s output (see Fig. 1).
3.1 Peer Review Task Formulation
To systematically evaluate the capabilities and robustness of LLMs in automated peer reviewing, we define two primary tasks that emulate essential components of the human review process: (1) Review Generation and (2) Score Prediction. These tasks are respectively designed to assess the LLM’s ability to produce comprehensive written feedback and assign quantitative evaluations across predefined criteria.
Task 1: Review Generation. The most direct application of leveraging LLMs as reviewers is the automatic generation of reviews for academic papers. In this task, we investigate the capability of LLM to produce constructive peer reviews that are coherent and consistent with human evaluations. Building on the aspect labels provided by Yuan et al. (2022), we construct a predefined aspect typology aligned with the ACL review guidelines. This typology includes the categories: “Summary”, “Motivation”, “Substance”, “Originality”, “Soundness”, “Clarity”, “Replicability”, “Meaningful comparison”, and “None”. Except for “Summary”, each aspect is further classified into positive and negative sentiments to capture both strengths and weaknesses of the paper. The “None” tag is assigned to sentences that do not fit any categories.
Using this typology, we prompt LLMs to write a review in the style of ICLR (International Conference on Learning Representations), instructing the model to tag each sentence with the corresponding aspect type at the beginning of the sequence. This approach ensures that the generated review addresses all relevant evaluation criteria and provides a structured assessment of the paper. An illustration of an automatically generated review with aspect tags is shown in Fig. 2.
Figure 2: The example of aspect-tagged review generated by GPT-4o-mini.
Task 2: Score Prediction. Score prediction serves as a vital complement to review generation by offering quantitative evaluations of a paper across multiple aspects Kang et al. (2018); Zhou et al. (2024). Its structured and numerical format provides clearer evaluation standards and enables more consistent comparisons. In this task, we assess the capability of LLMs to assign scores from different perspectives, simulating the nuanced judgments typically made by human reviewers.
Following the setup from Zhou et al. (2024), we prompt the LLM to predict scores for predefined aspects: “overall,” “substance,” “appropriateness,” “meaningful comparison,” “soundness correctness,” “originality,” “clarity,” and “impact.” Each score is expected to fall within a specific range (e.g., 1 to 10), in accordance with the scoring conventions of the target conference. This setup allows us to evaluate not only the LLM’s understanding of the paper but also its ability to apply evaluation criteria in a consistent and contextually appropriate manner.
To provide additional insight into the model’s reasoning, we ask it to include a brief justification for each score (as illustrated in Appendix C). This practice mirrors that of human reviewers, who often accompany their ratings with explanatory comments. By comparing the LLM-generated scores and justifications with those provided by human reviewers, we can assess the alignment between automated and expert evaluations. Furthermore, this task enables us to study the impact of adversarial text perturbations on the model’s scoring behavior. By introducing subtle modifications to the input, we can observe how the LLM’s assigned scores shift across different aspects. This analysis reveals the model’s sensitivity to input variation and its robustness in maintaining consistent evaluations under potentially manipulated conditions.
3.2 Attack Focus Localization
In paper reviewing tasks, the content under review is often too lengthy, making it computationally infeasible to directly search for adversarial patterns across the entire paper. To address this challenge, we first identify the content areas of the paper that attract the most attention to the LLM while generating reviews by feeding the clean paper content into the model and requesting a review.
Following this, we apply the Longest Common Subsequence (LCS) algorithm to identify matching subsequences between the paper content and the generated review. These matched subsequences are then highlighted as potential modifiable content for adversarial pattern injection. Let $x_{clean}$ represent the clean content of a single paper. For each paper, we extract a set of modifiable content by finding matching subsequences using LCS between the paper’s clean content and the review generated by the LLM. The process can be formalized as follows:
1. Identifying Modifiable Indexes. The modifiable subsequences are indexed by their start and end positions in the clean content:
$\mathcal{M}$ $=\text{LCS}\left(x_{\text{clean}},f_{r}(x_{clean})\right)$
=$=\{m_{1},m_{2},\dots,m_{N}\},$
where $\mathcal{M}$ is the modifiable sequences set for a single paper, $f_{r}(\cdot)$ denotes the review generation by LLM $f$ based on the input content. LCS identifies the longest common sequence matching algorithm.  [^1] LCS can be character-level or word-level depending on the attack method; implementation details could be found in Appendix D., $m$ represent modifiable subsequences, and $N$ is the total number of matched subsequences.
2. Searching Optimal Adversarial Pattern. The objective of these attacks is to find perturbations that lead to the maximum score shift for each paper, mathematically formulated as:
$\mathcal{P}_{m}=A\left(x_{clean},m\right)\ ,\forall m\in\mathcal{M}$
$X_{\text{adv}}=\left\{\left(x_{\text{clean}}\setminus m\right)\cup\left(m+\delta\right)\mid\forall m\in M,\forall\delta\in P_{m}\right\}$
$x_{adv}^{*}=\arg\max_{x\in\mathcal{X}_{adv}}\left(f_{s}(x_{adv}-f_{s}(x_{clean}))\right)$
, where $\mathcal{P}_{m}$ defines the possible set of perturbations for every modifiable sequence $m$, identified through the transformations specified by the attack method $A$. $\delta$ represents the textual perturbation applied to the original content, and $x^{*}_{adv}$ is the optimal perturbed content found through the search method of the attack. $f_{s}(\cdot)$ represents the sum of the score for all aspects predicted by LLM $f$.
The ultimate goal of the attack is to maximize the total score shift across all papers, expressed as:
Total Score $\text{Total Score Shift}=\sum^{|\mathcal{D}|}\left(f_{s}(x^{*}_{adv})-f_{s}(x_{clean})\right)$
where $\mathcal{D}$ represents the dataset of papers.
3.3 Textual Adversarial Attacks
Here, we introduce several textual adversarial attacks to explore the robustness of the review generation and score prediction procedure. We categorize adversarial attack methods based on common strategies in textual attacks. For each strategy, we introduce widely recognized representative approaches in black-box settings to test the limits of the robustness of the LLM.
Character-Level Attacks. For character-level adversarial attacks, we employ DeepWordBug Gao et al. (2018) and PuncAttack Formento et al. (2023). These methods mainly involve introducing typographical or grammatical errors in words by adding, deleting, repeating, replacing, or swapping characters or punctuation marks. Such perturbations can significantly affect the model’s performance while being less noticeable to humans.
Word-Level Attacks. For word-level adversarial attacks, we implement TextFooler Jin et al. (2020) and BERT-Attack Li et al. (2020) to deceive the LLM. Both attacks focus on replacing words with their synonyms or similar terms in contextualized embedding space, maintaining semantic similarity while altering the input text.
Sentence-Level Attacks. For the sentence-level attacks, we use StyleAdv Qi et al. (2021a) as representatives for sentence-level attacks. These methods involve transforming sentences in the paper content into specific text styles and modifying stylistic elements without altering the semantic meaning. Such transformation are designed to be imperceptible to human readers but can potentially mislead the LLM.
To conclude, applying these adversarial attack methods enables us to assess the robustness of LLMs in automated peer review. By introducing perturbations at the character, word, and sentence levels, we identify vulnerabilities that compromise the LLM’s reliability. The examples of each attack are provided in Table 5, which offers a visualizations of how these perturbations alter the original sentences. Also, the algorithm of textual adversarial attacks can be found in Appendix E.
4 Experiment
4.1 Experiment Settings
LLM Selection and Configuration. We adopt both closed- and open-source LLMs for automated peer review. From the closed-source family, we select GPT-4o and GPT-4o-mini due to their strong performance in text evaluation tasks Robertson (2023); Zhou et al. (2024), with the latter offering a cost-efficient alternative. For open-source models, we include LLaMA-3.3-70B and Mistral-small-3.1. Llama-3.3-70B is chosen for its strong general language capabilities, while Mistral-Small-3.1 has shown competitive performance with GPT-4o-mini, particularly in long-context understanding. Prompt templates and configurations are provided in Appendix A and B.
Dataset. PeerRead dataset Kang et al. (2018) is one of the earliest released datasets of peer reviews from the NLP community. Following Zhou et al. (2024), we select the ICLR-2017 subset for both review generation and score prediction tasks, which contains 427 papers (177 accepted and 255 rejected) with 1,304 reviews written by experts and aspect-specific scores assigned by the annotators. (The experiment result of Adversarial Attack Performance and Score Shift Comparison is provided in Appendix G)
AgentReview Jin et al. (2024) presents a large-scale dataset curated through their LLM-based peer review simulation framework. The dataset comprises papers from ICLR over four years (2020–2023), sourced from OpenReview, and includes 523 papers (350 rejected, 125 poster, 29 spotlight, and 19 oral) with 10,460 generated reviews on them.
Since the reviews in the both dataset are not manually annotated with aspect, we leverage the aspect tagger pre-trained on ASAP dataset Yuan et al. (2022) to automatically provide aspect tags for human-written reviews in the PeerRead dataset.
4.2 Evaluation Metrics
Review Quality Metrics. To address the question - How effective are LLMs in generating coherent and insightful reviews compared to human reviewers? We compare LLM-generated reviews with human-annotated reviews, treating the latter as ground truth. Following the metrics proposed by Yuan et al. (2022), we evaluate the text quality of the automated reviews across multiple dimensions. We aim to assess both the overall quality of the LLM-generated reviews and their consistency with human evaluations.
(1) Comprehensiveness: For automatically generated reviews, we expect them can cover diverse aspects of the paper. We calculate the Aspect Coverage (ACOV) metric to assess the generated review. In particular, the number of aspects in a predetermined aspect typology that a review has covered can be used to quantify aspect coverage.
$ACOV=\frac{\text{\# of aspect tags in generated review}}{\text{\# of predefined aspects}}$
(2) Similarity to human: To evaluate the semantic equivalence between LLM-generated reviews and multiple human-annotated ground truth reviews, we utilize classical statistical methods: ROUGE-1, ROUGE-2, ROUGE-L Lin (2004), and the BertScore metric Zhang et al. (2019). Rather than averaging scores across all human references, we select the maximum score between the LLM-generated review and the human-written ground truths. This approach assumes that closely aligning with every reference is unnecessary; instead, identifying the best match offers a more meaningful assessment of semantic equivalence.
Notably, bias and omissions may definitely exist in manually annotated reviews/scores. However, the human-written evaluation can still be the gold reference to a certain extent.
Robustness Metrics. To quantitatively assess the robustness of the LLM in the face of adversarial attacks, we define the following metrics.
(1) Sensitivity Analysis: Building on prior work in adversarial attacks, we adopt two straightforward metrics to evaluate the robustness of LLMs as reviewers: Attack Success Rate (ASR) and score shift. An attack is considered successful if the score shift between the clean output and the attacked output is at least +1.0 in the total score. ASR is defined as the proportion of such successful attacks, providing a clear indicator of the LLM’s vulnerability under adversarial conditions.
(2) Statistical Measures: We use the Wilcoxon signed-rank test Woolson (2005) as the statistical support. With the p-value set to 0.05, we pair the score that is predicted based on clean content and attacked content and test if the hypothesis that there exists a significant positive score shift after applying an adversarial attack holds true.
4.3 Evaluation – Review Quality
Table 1 shows the automatic evaluation results for the quality of reviews generated by LLMs. We begin by categorizing the human-written reviews into two types for comparison: (1) Within-Paper Pairs: review pairs from the same paper, representing the semantic similarity of reviews written for the same context. These pairs are generated using all possible combinations of reviews for each paper (e.g., $C^{n}_{2}$ combinations for $n$ reviews per paper), (2) Across-Paper Pairs: review pairs that are extracted from different papers, symbolizing the semantic similarity of reviews written for unrelated contents. We take all possible combinations of reviews across the entire dataset ($C^{n}_{2}$ combinations for N total human-written ground truth reviews) and exclude pairs that are included in the within-paper pairs to construct this set. Due to the significantly larger number of Across-Paper Pairs (2,774,190 pairs), we introduce random sampling to select 10% of the samples to create a subset (277,419 pairs), which is about 47 times larger than the 5,884 within-paper pairs. We observe that:
The ACOV of LLMs generated review is higher than human-written ground truths. The significance is further proved by Wilcoxon signed-rank test. The result can be attributed to the structured review guidelines provided in the prompt, which guide the model to cover a more comprehensive range of aspects. This suggests that LLMs have great potential to provide comprehensive perspectives on review principles in both positive and negative sentiments, demonstrating how LLMs can be guided to focus on desired qualities in reviews.
Within-paper pairs exhibit higher ROUGE scores. Reviews for the same paper consistently achieve higher R-1, R-2, and R-L scores, reflecting shared vocabulary, content, and structural similarity due to discussing the same paper-specific details. In contrast, across-paper pairs naturally show lower scores given their divergent content. LLM-generated reviews also show this trend: while slightly lower than human within-paper pairs, their ROUGE scores remain substantially higher than across-paper pairs, indicating the model’s ability to integrate paper-specific information and generate coherent and contextually aligned reviews.
Across-paper pairs exhibit BERTScore values comparable to within-paper pairs. This may be attributed to the semantic embeddings capturing shared linguistic structures or standardized grading language commonly used by reviewers, despite differences in paper content. For LLM-generated reviews, the similarity in BERTScore across both pair types indicates that LLMs are capable of generating reviews with structural and linguistic characteristics similar to those of human reviewers.
Table 1: Evaluation of LLM-generated review’s Comprehensiveness and Similarity to human. ACOV stands for Aspect Coverage. R-1, R-2, R-L, and BS indicate the Rouge-1/2/L, and BertScore, respectively. Human (within) reports the mean scores obtained from all pairs of human reviews written for the same paper, while Human (across) reports the mean over pairs taken from different papers.
Reviewer ACOV R-1 R-2 R-L BS
Human (within) 0.5285 0.4148 0.0973 0.1899 0.8403
Human (across) 0.5285 0.2188 0.0285 0.1154 0.8468
GPT-4o-mini 0.6489 0.3938 0.820 0.1875 0.8341
GPT-4o 0.9151 0.3469 0.0678 0.1676 0.8251
Llama-3.3-70B 0.6489 0.3432 0.0750 0.1760 0.8208
Mistral-small-3.1 0.9202 0.3541 0.0737 0.1658 0.8220
Human-based review quality evaluation. While automatic metrics assess surface-level qualities like fluency and similarity to human-written text, they miss deeper aspects such as coherence, usefulness, and critical insight. To address this gap, we conduct a controlled human evaluation to complement the automatic analysis, as detailed in Appendix H.
4.4 Evaluation – Attack Effectiveness
Table 2: Adversarial Attack Performance and Score Shift Comparison of AgentReview Jin et al. (2024) dataset. “ASR” represents the Attack Success Rate. “Avg. Score shift” signifies the average score shift between attacked outputs and clean outputs. “Avg. #Pos/Neg Shift” indicates the average change in the number of positive and negative tags assigned to the reviews. “#Queries” refers to the average number of querying LLM. The value are reported as result from GPT-4o-mini / GPT-4o / Llama-3.3-70B / Mistral-Small-3.1.
Attacks ASR $\uparrow$ Average Shift #Queries
Score $\uparrow$ #Pos Tags $\uparrow$ #Neg Tags $\downarrow$
DeepWordBug 0.87 / 0.90 / 0.81 / 0.91 3.44 / 4.41 / 2.56 / 3.55 0.24 / 0.01 / -0.08 / -0.02 0.78 / 1.49 / 1.70 / 0.32 43 / 51 / 52 / 55
PuncAttack 0.85 / 0.87 / 0.77 / 0.87 3.44 / 4.05 / 2.52 / 3.59 0.27 / 0.01 / -0.11 / -0.01 0.82 / 1.12 / 1.79 / 0.19 51 / 64 / 66 / 69
TextFooler 0.89 / 0.92 / 0.85 / 0.91 3.77 / 4.71 / 2.98 / 3.94 0.33 / 0.00 / -0.15 / -0.03 0.84 / 1.23 / 2.11 / 0.37 47 / 52 / 58 / 58
BERT-Attack 0.87 / 0.89 / 0.71 / 0.90 3.59 / 4.65 / 2.36 / 4.17 0.36 / -0.01 / -0.06 / -0.03 0.69 / 1.52 / 1.51 / 0.47 54 / 56 / 56 / 66
StyleAdv 0.89 / 0.92 / 0.82 / 0.92 3.62 / 4.73 / 2.87 / 3.97 0.30 / 0.01 / -0.16 / -0.04 0.82 / 1.31 / 2.25 / 0.38 25 / 32 / 25 / 30
As discussed in Section 3.3, we investigate the robustness of LLM-based reviewers by applying widely recognized textual adversarial attacks to manipulate input paper content. Table 2 summarizes the results of these attack baselines aimed at artificially inflating the review scores predicted by LLMs. We observe that:
Most attacks achieve high attack success rate (ASR). ASR among different granualarities of attacks ranging from 71% to 92%, accompanied by substantial score shifts. These findings demonstrate the effectiveness of adversarial methods in deceiving LLM-based reviewers and highlight the models’ sensitivity to such subtle manipulations injected into a lengthy document.
Changes in the distribution of review sentiment tags. Specifically, the average increase in positive tags (Avg. #Pos Shift) correlates with higher positive sentiment scores after attacks, while the average number of negative tags (Avg. #Neg Shift) consistently decreases. This suggests that even minor content perturbations can bias LLMs toward generating more favorable feedback, suppressing critical evaluations. As the predicted scores inflate, the models tend to produce more positive assessments while reducing the expression of negative ones, leading to overly optimistic reviews that lack the balanced critiques typical of human reviewers.
Negative tag suppression aligns with inflated scores. Although positive tags sometimes decrease after attacks, the reduction in negative tags is generally more significant. This trend is particularly evident in the Llama-3.3-70B and Mistral-small-3.1, where adversarial modifications substantially reduce the number of negative aspect tags. One possible explanation is that these models inherently favor positive or neutral sentiment generation, making them more susceptible to adversarial manipulations that suppress negative feedback. This imbalance suggests that adversarial attacks effectively weaken the models’ capacity to generate critical-depth reviews that are supposed to be provided in the peer-reviewing process.
To statistically validate the impact of adversarial attacks, we conducted a Wilcoxon signed-rank test to assess whether post-attack scores significantly exceed pre-attack scores. As shown in Table 2, the average score shifts are statistically significant, with p-values consistently below 0.05. These results confirm that the observed score inflation is not due to random variation but is a direct consequence of the applied adversarial manipulations.
Fig. 3 visualizes the average scores across various aspects in the predefined typology, comparing the results before and after applying StyleAdv to the input paper content. The radar chart shows that post-attack scores generally surpass pre-attack scores across most aspects, highlighting the attack’s effectiveness. Notably, some aspects, such as originality, clarity, and appropriateness, appear less sensitive to stylistic changes, likely because modifying writing style alone does not significantly alter the semantic meaning of the content. On the other hand, aspects like substance and meaningful comparison show greater vulnerability, as stylistic manipulation may reduce the depth of supporting evidence or coherence in comparisons, thus impacting the prediction more easily.
Figure 3: visualization of the average score of each aspect predicted by GPT-4o-mini. Clean denotes the prediction when input is not manipulated by StyleAdv.
Human-based attack imperceptibility evaluation. In addition to the quantitative results in Table 2, we conducted a user study to assess the stealthiness of attacks in a realistic peer review setting. Specifically, we evaluated participants’ ability to identify attacked papers and localize the perturbed paragraphs. The detailed settings and full user study results are reported in Appendix H.
5 Limitations of Possible Defenses
While recent studies have proposed using AI-generated text detectors (e.g., GPTZero, OpenAI Classifier) to identify LLM-authored reviews or adversarially manipulated inputs, such detection-based defenses face fundamental limitations when confronting the vulnerabilities in the peer-review process revealed in this work.
Detecting Adversarial Manipulations in the Paper Content.
To apply detectors like GPTZero directly to the paper content to identify whether it has been adversarially manipulated (e.g., injected with typos or stylistic changes to mislead LLM reviewers) is beyond the current capabilities of AI-text detectors. These detectors are typically trained to distinguish between human-written and LLM-generated text, rather than to identify subtle, semantics-preserving perturbations designed to mislead LLM reviewers. These perturbations are designed to evade both human perception and statistical detectors, as confirmed by our human evaluation results. Thus, applying existing detectors to the paper content is unlikely to surface such attacks, and developing dedicated perturbation-aware detectors remains an important future direction.
Detecting Machine-generated Reviews.
A plausible use of detectors is to identify LLM-generated reviews, with the goal of adjusting confidence scores or triggering human audits. However, this strategy suffers from several challenges.
Policy Ambiguity. Due to current conference policies (e.g., the ACL Policy on Publication Ethics [^2] https://www.aclweb.org/adminwiki/index.php/ACL_Policy_on_Publication_Ethics#Guidelines_for_Generative_Assistance_in_Authorship and the NeurIPS 2025 LLM Policy [^3] https://neurips.cc/Conferences/2025/LLM) explicitly allowing the use of LLMs for language refinement, there is no clear boundary between LLM-written and LLM-assisted reviews. Applying a strict detection threshold risks unfairly penalizing legitimate human-written reviews that simply use LLMs for improving wording and fluency.
Ease of Bypassing Detection. Existing detectors are vulnerable under various real-world application scenarios. An adversary can paraphrase LLM-generated reviews Krishna et al. (2023) or design prompts to guide LLMs in generating human-like text to bypass detectors Wu et al. (2024), while these reviews still rely on LLM content that lacks a deep understanding of the paper. This undermines the goal of ensuring high-quality reviews. Futhermore, as conferences begin to officially adopt LLM tools in the review pipeline (e.g., AAAI’s AI-powered peer review assessment system [^4] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/), relying solely on detector-based filtering appears inadequate to address the vulnerabilities introduced by adversarial manipulations of the paper content.
These challenges underscore the vulnerability of LLM-assisted review systems. The adversarial attacks are not only effective but also difficult to defend against using current methods.
6 Conclusion
We introduce a framework for evaluating the capabilities and robustness of large language models (LLMs) in automated peer-review tasks, addressing the significant burden on reviewers from the surge in manuscript submissions. Utilizing mainstream LLMs, we perform two core tasks—review generation and score prediction—demonstrating the effectiveness of LLMs in providing coherent and insightful feedback. To assess robustness, we apply various adversarial techniques to manipulate paper content and observe their impact on the LLM’s predicted scores, revealing vulnerabilities where subtle input changes significantly affect reliability. In response, we plan to develop defense mechanisms, such as inference-time detection methods like perplexity detection, to enhance the stability of LLM-based peer review processes.
7 Limitation
Despite the valuable insights provided by this study, several limitations should be acknowledged. First, the evaluation of the LLM’s performance was limited to the chosen representative models (i.e., GPT-4o, GPT-4o-mini, Llama-3.3-70B, and Mistral-Small-3.1), and the generalizability of the findings to other LLMs remains uncertain. Different models may exhibit varying levels of robustness and accuracy, and future work should explore a broader range of models. Additionally, the scope of textual adversarial attacks applied in this research was restricted to certain common techniques. More sophisticated or domain-specific adversarial methods (e.g., adaptive attacks, non-black-box attacks) could further challenge the LLM’s performance, which was not considered in this study. Finally, the focus on review generation and score prediction leaves out other critical aspects of the peer review process, such as the interaction between reviewers and authors or the integration of LLM-generated reviews into existing peer review workflows, which warrant further investigation.
8 Ethical Considerations
This work does not rationalize or encourage the use of LLMs in the paper reviewing process. Instead, our aim is to provide a comprehensive discussion on the generation patterns of LLM-based reviewers and raise awareness of the vulnerabilities in applying LLM-generated results, as they can be easily manipulated through simple textual attacks and may lack the expertise and proper judgment required for such tasks.
Bias and Fairness.  LLMs are prone to biases based on their training data and also potentially suffer from position, verbosity, and self-enhancement biases, which may affect their fairness in paper review tasks. Ethical guidelines suggest conducting rigorous testing and mitigation strategies to reduce bias, ensuring that the reviews or scores generated are not distorted by manipulations in textual content that are introduced intentionally or not. Human reviewers should remain responsible for the final review and decision, even if LLMs assist in the process. Ethical AI usage involves maintaining human oversight to ensure accountability and avoid transferring full decision-making power to an algorithm that may lack context or moral judgment. AI Usage Compliance.  Our research testing the robustness of LLM reviewers to adversarial attacks should adhere to ethical standards of security and integrity. This involves ensuring that such testing is conducted in a controlled environment, with clear goals of improving AI reliability rather than exploiting vulnerabilities for harmful purposes. Such experiments aim to strengthen the resilience and trustworthiness of AI systems.
Fragility of AI-text Detectors.  While utilizing AI-text detectors to identify machine-generated content appears to be a potential defense against our proposed threats, such mechanisms are insufficient to ensure security in real-world peer reviewing scenario for several reasons (detailed in Section 5).
First, current detectors are primarily trained to distinguish between human-written and LLM-generated text, but they lack the capability to detect subtle, semantics-preserving adversarial perturbations in paper content. Second, zero-shot detection of LLM-generated reviews can be trivially bypassed through paraphrasing or minor editing, allowing AI-text to evade filtering. Moereover, the increasing integration of LLMs into legitimate writing and reviewing workflows—often under ambiguous policy guidelines—renders binary detections ethically problematic and operationally difficult.
From an ethical standpoint, it is essential to critically assess that blanket reliance on such tools may foster false confidence while unintentionally marginalizing non-native English speakers who use LLMs for language assistance. While AI-text detectors may serve as a useful complementary measure, they are not by themselves a sufficient defense against the challenges we uncover.
In all cases, human oversight remains essential for ensuring the fairness, accountability, and integrity of peer review.
Appendix A Prompt Example
Fig. 4 presents the detailed prompt template used for automatic review generation in our experiments. The prompt is explicitly crafted to simulate the role of a professional reviewer specialized in computer science and machine learning, guiding LLMs to produce structured reviews in ICLR style. To ensure comprehensive critique, the prompt directs the model to include both positive and negative aspects of the paper by emphasizing: “Make sure to include both positive and negative aspects.” This approach addresses findings from Zhou et al. (2024), who noted that LLMs often generate overly positive reviews, risking optimistic evaluations that overlook a paper’s weaknesses. By encouraging balanced feedback, the prompt aims to produce more critical and informative assessments.
The prompt further enhances review quality by tagging sentences with specific aspects, ensuring fine-grained analysis and differentiation of review components. It also requires the model to assign scores across multiple evaluation dimensions, with brief explanations for each, enabling quantitative assessment aligned with qualitative commentary. Additionally, to maintain conciseness and consistency, we specified that the total output should not exceed 500 tokens. This design promotes consistent, reproducible review generation that mirrors real-world peer review processes, allowing us to systematically evaluate the robustness and interpretability of LLM-based reviewing systems.
Prompt Template You are a professional reviewer in computer science and machine learning. Based on the given content of a research paper, you need to write a review in ICLR style and tag sentences with the corresponding tag type at the beginning of sequence: tags types: [NONE], [SUMMARY], [MOTIVATION POSITIVE], [MOTIVATION NEGATIVE], [SUBSTANCE POSITIVE], [SUBSTANCE NEGATIVE], [ORIGINALITY POSITIVE], [ORIGINALITY NEGATIVE], [SOUNDNESS POSITIVE], [SOUNDNESS NEGATIVE], [CLARITY POSITIVE], [CLARITY NEGATIVE], [REPLICABILITY POSITIVE], [REPLICABILITY NEGATIVE], [MEANINGFUL COMPARISON POSITIVE], [MEANINGFUL COMPARISON NEGATIVE]. Your total output should not surpass 500 tokens, make sure to include both positive and negative aspects. Also, you need to predict the review score in several aspects based on the generated review, providing an explanation of each aspect in less than 30 tokens. Choose a integer score from 1 to 10, higher score means better paper quality. Section {section name 1}: {section content 1} Section {section name 2}: {section content 2} Section {section name N}: {section content N} Please strictly follow the format of Example output: 1. REVIEW: tagged sequences. 2. REVIEW SCORE: OVERALL: score, SUBSTANCE: score, APPROPRIATENESS: score, MEANINGFUL COMPARISON: score, SOUNDNESS CORRECTNESS: score, ORIGINALITY: score, CLARITY: score, IMPACT: score. 3. REVIEW SCORE EXPLANATION: OVERALL: explanation, SUBSTANCE: explanation, APPROPRIATENESS: explanation, MEANINGFUL COMPARISON: explanation, SOUNDNESS CORRECTNESS: explanation, ORIGINALITY: explanation, CLARITY: explanation, IMPACT: explanation.
Figure 4: Prompt template used for review generation.
Appendix B Implementation Details – Model Configuration
We accessed closed-source models (GPT-4o and GPT-4o-mini) via the OpenAI API, and open-source models—Llama-3.3-70B (approximately 70 billion parameters, running on H100 via vLLM, requiring around 140GB of memory) and Mistral-Small-3.1 (approximately 24 billion parameters, running on H100 via vLLM, requiring around 60GB)—through Hugging Face. All models were configured with a temperature of 0.3 to balance creativity and coherence. The maximum number of generated tokens was set to 2048, ensuring detailed reviews without exceeding practical length constraints.
Appendix C Case Study
Table 3 presents aspect-wise scores predicted by GPT-4o-mini along with corresponding explanations. With access to the full paper content, the model generates fluent and coherent justifications that highlight both strengths (Substance, Soundness, Originality) and areas for improvement (Clarity, Appropriateness). These explanations enhance the interpretability of the scoring process and demonstrate the model’s ability to reason about multiple dimensions of scholarly quality. Despite full access to the paper, the predicted scores still cluster within a relatively narrow range (6–8), indicating a limited critical depth. This suggests that LLM may default to generally favorable assessments unless explicitly prompted to be more discriminative. Overall, the generated explanations not only offer transparency into the model’s evaluative reasoning but also reveal opportunities for further alignment with human reviewing norms.
Table 3: Aspect score predicted by GPT-4o-mini with the explanations.
Aspect Score Explanation
Overall 7.0 The paper makes significant contributions to GAN theory but has areas for improvement in clarity and practical implications.
Substance 8.0 The rigorous theoretical analysis and proofs are strong, providing a solid foundation for the claims.
Appropriateness 6.0 The motivation is relevant, but the practical implications could be better articulated for broader impact.
Meaningful Comparison 7.0 The paper contextualizes its contributions well but could engage more critically with alternative approaches.
Soundness Correctness 8.0 The theoretical framework is robust, though some assumptions may limit applicability.
Originality 7.0 The introduction of new tools is commendable, but the paper could better distinguish its contributions from existing work.
Clarity 6.0 While generally clear, some sections are dense and could be simplified for accessibility.
Impact 7.0 The findings have the potential to influence future research directions in GAN training, though practical applications need clearer emphasis.
Appendix D Implementation Details – Longgest Common Sequence (LCS) split
The granularity of modifiable content found by the LCS algorithm is dynamic, as we split both the clean paper content and the generated review into segments based on the chosen attack level, which can be character-level, word-level, or sentence-level. The LCS matching is then performed at the appropriate granularity.
Character-level Attack: The LCS algorithm then identifies the longest common subsequences at the character fine-grained level. This allows for precise modifications at the character level, which is useful for attacks that require small, detailed changes.
Word-level Attack: For a word-level attack, both the clean paper content and the generated review are split by individual words. The LCS matching process is constrained to operate on these word-level splits, as word-level attacks typically involve replacing words with synonyms or other word-based modifications. This approach ensures that each modifiable subsequence corresponds to complete words, enabling word-level changes to occur seamlessly.
Sentence-level Attack: For sentence-level attacks, the LCS is first performed at the word level, identifying the longest common subsequences between words in the clean content and the review. Once the LCS matches are found at the word level, the matched sequence is expanded to include the full sentence. Specifically, for any word sequence that forms a matched subsequence, we extend the match forward and backward to capture the entire sentence in which the sequence appears. This ensures that the adversarial patterns can be injected into complete sentences, preserving the overall syntactic and semantic structure of the content.
Appendix E Implementation Details – Textual Adversarial Attack.
Textual attacks are composed of four fundamental components: (1) Goal Function defines the objective of the attack, offering a heuristic for search strategies to identify optimal solutions. It evaluates how effectively perturbations achieve desired outcomes, such as convincing LLMs to favor the altered paper for better ratings and feedback. (2) Constraints ensure that the disturbed text maintains a resemblance to the original, preserving semantic similarity to validate the modification. (3) Transformation process details methods for altering input content, including character modifications, synonym replacements, and changes in style and syntax. (4) Search Method explores valid perturbations produced by transformations, employing techniques such as greedy search with word importance ranking, beam search, and brute-force search.
The textual attack methods we used to manipulate the paper content follow the aforementioned formulation and are primarily implemented using TextAttack [^5] https://github.com/QData/TextAttack Morris et al. (2020). Specifically, the goal function for all the employed attack methods-namely, DeepWordBug, PuncAttack, TextFooler, BERT-Attack, and StyleAdv is to maximize the total score shift. For the constraint, we basically follow the setting in PromptBench  Zhu et al. (2023) and set the max candidate (maximum number of generated perturbation candidates) of TextFooler and BERT-Attack to 15.
For the search part, we adopt a greedy approach with word importance ranking for DeepWordBug, BERT-Attack, PuncAttack, and TextFooler. We restrict our process to the top 50 most important words to balance computational affordability and attack effectiveness, feeding these into our perturbation generation pipeline. StyleAdv uses brute-force at the sentence level because both methods do not involve in word-level searching. We segment paper content into individual sentences using the NLTK tokenizer [^6] https://www.nltk.org/api/nltk.tokenize.html Bird (2006). For style attacks, we utilize the pre-trained model Krishna et al. (2020) to convert text into Bible style. Original sentences are replaced with their transformed counterparts as adversarial examples.
Appendix F Degree of changes using different adversarial attacks
To further quantify the degree of changes when applying different textual adversarial attacks on paper content, we report two standard metrics Modification Rate and Semantic Similarity.
Given a pair of paper content—original $x_{\text{clean}}$ and adversarial $x_{\text{adv}}$—we first extract the subset of sentences that differ using sentence-level alignment and comparison. This results in a filtered set of sentence pairs:
$\mathcal{S}=\big{\{}(s_{\text{clean}},s_{\text{adv}})\ \big{|}$ $s_{\text{clean}}\in x_{\text{clean}},s_{\text{adv}}\in x_{\text{adv}},$
$\ s_{\text{clean}}\neq s_{\text{adv}}\big{\}}$
Each pair $(s_{\text{clean}},s_{\text{adv}})\in\mathcal{S}$ contains a sentence from the original and its corresponding modified version from the adversarial input. We then compute two metrics over this set.
  • Modification Rate, which measures the extent of perturbation introduced into the original text.
    For each sentence pair in $(s_{\text{clean}},s_{\text{adv}})\in\mathcal{S}$ , the normalized character-level change is:
    Modification $\text{Modification Rate}=\frac{\Delta_{\text{char}}(s_{\text{clean}},s_{\text{adv}})}{(|s_{\text{clean}}|+|s_{\text{adv}}|)/2}$
    where $\Delta_{\text{char}}(s_{\text{clean}},s_{\text{adv}})$ denotes the number of changed characters (insertions, deletions, or substitutions).
  • Semantic Similarity, measured by BERTScore, to assess how closely the adversarial texts $s_{\text{adv}}$ preserve the original semantics $s_{\text{clean}}$
Our results in Table 4 show that character-level attacks (e.g., DeepWordBug, PuncAttack) introduce minimal changes, with modification rates below 1% and BERTScore exceeding 0.97, indicating near-perfect semantic preservation. Word-level attacks such as TextFooler and BERT-Attack introduce slightly higher modification rates (around 7%–14%) while still maintaining high semantic similarity (above 0.98). In contrast, the sentence-level attack StyleAdv results in a substantially higher modification rate (over 130%) and a more noticeable drop in semantic similarity (BERTScore of 0.8558). This is expected, as StyleAdv explicitly targets writing style transformation, which involves rephrasing and substituting multiple words throughout the sentence. As a result, even though the underlying meaning is largely preserved, the surface form diverges significantly from the original. These findings demonstrate that most adversarial modifications are subtle and semantically faithful—except for sentence-level style attacks, which are more aggressive in altering surface text while still maintaining reasonable semantic alignment. This confirms that the adversarial examples are both realistic and challenging to detect.
Table 4: Quantitative comparison of text changes under different attack methods. We report both character-level Modification Rate and Semantic Similarity (BERTScore).
Attack Type Modification Rate $\downarrow$ Semantic Similarity $\uparrow$
DeepWordBug Char 0.0090 0.9816
PuncAttack Char 0.0073 0.9749
TextFooler Word 0.0761 0.9885
BERT-Attack Word 0.1362 0.9828
StyleAdv Sent 1.3117 0.8558
In addition, to qualitatively evaluate the degree of changes, we have provided more examples in Table 5 shown below.
From these examples in the table, we can find that character-level attacks like DeepWordBug and PuncAttack make minimal, often imperceptible changes that nonetheless disrupt model predictions, revealing vulnerability to small spelling errors. Word-level attacks (TextFooler, BERT-Attack) replace words with adversarial synonyms, maintaining syntactic structure and semantics but misleading models reliant on lexical cues. Sentence-level attacks (StyleAdv) introduce broader stylistic and phrasing changes, preserving overall meaning but posing challenges for detection. These results highlight a trade-off between semantic preservation and attack stealthiness, emphasizing the need for defenses that handle perturbations from character edits to stylistic rewrites.
Table 5: Examples of content under different levels of textual adversarial attacks. The Type column indicates the attack granularity (character-, word-, or sentence-level), and the injected perturbations are highlighted in red.
Attack Type Example Sentence
Clean - The model can also be used in conjunction with differentiable memory mechanisms for implicit relation discovery in one-shot learning tasks.
DeepWordBug Char The model can also be used in conjufnction with differentiable memory mechanisms for implicit relation discovery in one-shot learning tasks.
PuncAttack Char The model can also be used in c&onjunction with differentiable memory mechanisms for implicit relation discovery in one-shot learning tasks.
TextFooler Word The paragon can also be used in conjunction with differentiable memory mechanisms for implicit relation discovery in one-shot learning tasks.
BERT-Attack Word The model can also be used in conjunction with differentiable memory mechanisms for explicit relation discovery in one-shot learning tasks.
StyleAdv Sent The model can also be used in conjunction with a variety of differentiable memory mechanisms for the implicit relation discovery in one shot learning tasks.
Appendix G Evaluation – Adversarial Attack Result for PeerRead Dataset
Table 6: Adversarial Attack Performance and Score Shift Comparison of PeerRead Kang et al. (2018) dataset. “ASR” represents the Attack Success Rate. “Avg. Score shift” signifies the average score shift between attacked outputs and clean outputs. “Avg. #Pos/Neg Shift” indicates the average change in the number of positive and negative tags assigned to the reviews. “#Queries” refers to the average number of querying LLM. The value are reported as result from GPT-4o-mini / GPT-4o/ Llama-3.3-70B / Mistral-Small-3.1.
Attacks ASR $\uparrow$ Average Shift #Queries
Score $\uparrow$ #Pos Tags $\uparrow$ #Neg Tags $\downarrow$
DeepWordBug 0.92 / 0.81 / 0.82 / 0.89 4.27 / 3.87 / 3.06 / 3.22 0.33 / 0.00 / 0.07 / -0.02 -0.95 / 1.18 / -1.73 / -0.24 42 / 51 / 50 / 56
PuncAttack 0.87 / 0.90 / 0.79 / 0.83 3.90 / 4.23 / 3.17 / 2.98 0.33 / -0.03 / 0.03 / -0.02 -0.81 / 1.18 / -1.71 / -0.20 51 / 63 / 64 / 68
TextFooler 0.92 / 0.95 / 0.87 / 0.91 4.61 / 4.30 / 2.87 / 3.44 0.35 / -0.03 / -0.07 / -0.01 -0.90 / 1.50 / -2.14 / -0.32 46 / 54 / 57 / 58
BERT-Attack 0.85 / 0.74 / 0.70 / 0.85 4.13 / 4.43 / 2.35 / 3.64 0.28 / 0.03/ -0.09 / -0.02 -0.75 / 1.23 / -1.70 / -0.40 52 / 53 / 54 / 65
StyleAdv 0.83 / 0.88 / 0.80 / 0.87 3.58 / 4.28 / 2.35 / 2.81 0.27 / -0.10 / -0.14 / -0.02 -0.65 / 1.13 / -1.96 / -0.26 12 / 16 / 12 / 16
Table 6 presents the results of different attack methods aimed at intentionally boosting the scores assigned by the LLM-based reviewer. Based on the results, we draw the same trend as shown in Table 2. This demonstrates the effectiveness of the attack across different datasets and further confirms the vulnerability of LLMs to simple textual adversarial attacks under a black-box scenario.
A closer analysis of the results reveals that:
(1) TextFooler consistently achieves the highest Attack Success Rate (ASR) across models, with a peak of 0.95 for GPT-4o, suggesting that even minor lexical substitutions can significantly distort LLMs’ judgment. This high ASR is paired with a substantial average score shift, particularly for GPT-4o-mini, indicating that such perturbations can systematically induce inflated evaluations. The consistent increase in positive tags and reduction in negative tags further highlights a dual manipulation effect—enhancing positive sentiment while suppressing critical feedback—directly compromising review integrity, aligning with the increased predicted score.
(2) StyleAdv, while maintaining a moderate ASR, achieves comparable score shifts with the lowest query count, making it the most efficient attack among the tested methods. This suggests that stylistic manipulations, even with minimal transformation candidates and model interaction, are particularly deceptive to LLMs, likely because of their characteristic in high modification rate and thus have space to alter the distribution of attacked content to differentiate from the original one.
(3) GPT-4o and GPT-4o-mini exhibit the highest susceptibility to adversarial attacks, Llama-3.3-70B and Mistral-Small-3.1 show relatively better resilience. We hypothesize this difference arises from GPT models’ stronger emphasis on maintaining contextual coherence. This strength in natural text generation makes them more prone to retaining even sentiment shifts caused by manipulations as coherent opinions, leading to more pronounced score shifts. In contrast, Llama-3.3-70B’s architecture may prioritize semantic understanding over coherence, making it slightly more robust against such manipulations.
These findings emphasize the critical need for developing more robust LLMs, especially in high-stakes applications such as automated review generation. Beyond improving model architectures against existing textual adversarial attacks, future research should explore more resilient prompting techniques and processing mechanisms that can mitigate the impact of adversarial perturbations.
Appendix H Evaluation – Human Evaluation
We conduct a controlled human evaluation to complement automatic metrics and assess the quality and robustness of LLM-generated peer reviews. This evaluation is structured around two key research questions:
(1) How similar are LLM-generated reviews to human-written ones in terms of perceived quality?
(2) To what extent adversarial textual patterns of varying granularity can remain stealthy when embedded in paper content.
To answer these questions, we designed two complementary human studies: one focused on review quality evaluation, and another on attack stealthiness and identifiability. We collected user study results from 31 participants, all of whom are master’s or PhD students with background knowledge in deep learning and natural language processing. The setup and methodology is described below for each in turn.
H.1 Review Quality Ranking
While automatic evaluation metrics offer a broad assessment of LLM-generated reviews, demonstrating their ability to cover aspects and exhibit high semantic similarity to human-written reviews, they fall short in capturing the qualities of review usefulness, coherence, and perceived value from a human perspective. Specifically, such metrics may favor reviews that are grammatically fluent and lexically aligned with human-written text, even when such reviews lack substantive critique or meaningful engagement with the paper. To mitigate these limitations, we conduct a controlled human evaluation to complement automatic analysis and offer a more comprehensive understanding of review quality.
To evaluate human-perceived review quality, we adopt the following three criteria:
(1) Fluency & Consistency: Is the review easy to read and internally coherent?
(2) Content Relevance: Does the review engage with the paper’s core contributions, and does it possess alignment with the meta-review context?
(3) Insightfulness: Does the review provide substantive, constructive feedback that could help improve the paper?
For each review set, participants were asked to rank the four anonymized reviews (one LLM-generated and three human-written) along the above three quality dimensions. Following each ranking task, they also provided a confidence score (on a 1–5 scale) indicating the certainties about their judgments.
Result & Discussion. The human evaluation results in Table 7 reveal insights emerging regarding the human-perceived quality of LLM-generated reviews as follows:
(1) Lamas-3.3-70B consistently outperformed other models in all three ranking criteria, achieving the highest ratings (lowest Rank value) and the highest T-1 rate for Content Relevance, suggesting that it produces reviews most closely aligned with the core contributions of the evaluated articles. This superior performance aligns with the hypothesis that larger models are more proficient in grasping complex content.
(2) GPT-4o displayed a relatively weak performance in Insightfulness compared to other LLMs, with the lowest T-1 Rate and the highest average rank, indicating that it struggles to offer constructive, in-depth feedback, a critical factor in effective peer review. A potential explanation lies in GPT-4o’s strict adherence to our structured prompt, which encouraged balanced coverage of all review aspects but may have inadvertently constrained its ability to deliver focused, in-depth analysis of specific aspects. Notably, despite this limitation, GPT-4o’s reviews still outperformed at least one human-written review in the ranking tasks, highlighting its overall competitiveness.
(3) Mistral-small-3.1, despite demonstrating strong Fluency & Consistency with the highest T-1 Rate, lagged in Insightfulness, suggesting that fluency alone does not guarantee meaningful critique.
(4)GPT-4o-mini demonstrated a balanced performance, achieving competitive results across all criteria, especially in Insightfulness, indicating that smaller, more efficient models can still achieve strong human-perceived quality.
These results collectively highlight the importance of balancing fluency, relevance, and depth in LLM-generated reviews, with Llama-3.3-70B setting a strong benchmark in this regard. However, the somewhat underwhelming performance of models like GPT-4o in providing insightful feedback underscore the ongoing challenge of enhancing LLMs to generate not just coherent text but also substantive and contextually valuable content.
Table 7: User Study on Review Quality Ranking. “Rank” denotes the average position of the LLM-generated review among the three human reviews (lower is better), “T-1 Rate” is the proportion of times the LLM review was ranked first for that criterion. “Conf.” is the average confidence score users assigned to their rankings.
Attacks Fluency & Consistency Content Relevance Insightfulness
Rank $\downarrow$ T-1 Rate $\uparrow$ Conf. $\uparrow$ Rank $\downarrow$ T-1 Rate $\uparrow$ Conf. $\uparrow$ Rank $\downarrow$ T-1 Rate $\uparrow$ Conf. $\uparrow$
GPT-4o-mini 2.07 0.45 4.03 1.83 0.52 4.03 1.66 0.55 4.03
GPT-4o 2.31 0.31 3.93 2.31 0.38 3.93 2.66 0.17 3.93
Llama-3.3-70B 1.90 0.45 3.97 1.69 0.59 3.97 1.72 0.52 3.97
Mistral-small-3.1 2.10 0.52 3.72 1.93 0.41 3.72 2.41 0.24 3.72
H.2 Human-based Attack Imperceptibility Evaluation
To further examine the robustness of LLM-generated content, we conducted a human evaluation targeting the identifiability of adversarial perturbations under realistic review scenarios. Specifically, this experiment aimed to answer two key research questions:
(1) Attack Identification: In a simulated peer-review setting, can human evaluators distinguish which papers has been adversarially attacked?
(2) Attack Localization: When explicitly informed that the paper content has undergone an adversarial attack, can evaluators locate the specific paragraph containing the adversarial perturbation?
Settings. To this end, we recruited annotators with experience in academic writing to participate in a two-stage evaluation task. In the first stage, each annotator was presented with a pair of papers, one clean and one adversarially perturbed, and asked to identify which appeared adversarially manipulated. In the second stage, annotators were shown a single adversarially perturbed article and instructed to highlight the paragraph they believed was most indicative of adversarial editing via a given specific textual attack method.
Table 8: User Study of Attack Imperceptibility. The study consists of two parts: Attack Identification and Attack Localization. “Acc.” represents the accuracy of identifying the attacked paper or position, “Unc.” denotes the Uncertainty Rate, calculated as the proportion of times users responded with “N/A, can’t distinguish.” “Conf.” refers to the average confidence score users assigned to their answers.
Attacks Identification Localization
Acc. $\downarrow$ Unc. $\uparrow$ Conf. $\downarrow$ Acc. $\downarrow$ Unc. $\uparrow$ Conf. $\downarrow$
DeepWordBug 0.62 0.55 3.28 0.96 0.14 4.31
PuncAttack 0.88 0.41 4.10 0.69 0.10 4.48
TextFooler 0.63 0.45 3.45 0.26 0.07 3.17
BERT-Attack 0.50 0.72 3.48 0.08 0.10 2.86
StyleAdv 0.73 0.24 3.76 0.21 0.34 2.79
This setting reflects real-world reviewing conditions where reviewers rely solely on textual cues. The evaluation helps assess not only the perceptibility of adversarial patterns to informed readers but also whether these attacked papers can effectively evade human detection even under heightened scrutiny, highlighting the potential risk of such subtle attacks in influencing LLM-based review systems.
Result & Discussion. According to Table 8, we observe substantial variation in the imperceptibility of different granularities of adversarial attacks and discuss them as follows:
(1) BERT-Attack emerges as the most stealthy method, with annotators achieving only 50% accuracy in identifying perturbed papers (i.e., chance level), and a low localization accuracy of 8%. This suggests that semantic-preserving synonym substitution strategies are particularly effective in preserving the naturalness of text under human scrutiny, making them difficult to detect even when readers are explicitly informed of adversarial presence.
(2) Despite being originally designed to subtly change writing style, StyleAdv showed a relatively high identification accuracy (73%) but notably poor localization performance (21%), accompanied by the highest uncertainty rate (34%). This outcome suggests that stylometric perturbations may not blend well in academic contexts, possibly because the injected styles (i.e., biblical) deviate too far from the expected scholarly tone.
As detailed in Appendix F in our revision, StyleAdv also involves higher modification rates and lower semantic similarity, which may disrupt paragraph-level cohesion and topical consistency, making the manipulated version stand out when compared side-by-side with a clean reference. However, the poor localization accuracy indicates that in zero-shot settings, where no clean version is available for comparison, humans struggle to pinpoint the attack, highlighting the stealthy nature and potential risk of such stylistic perturbations. This threat may be further amplified if future attacks adopt writing styles that are more plausible within academic domains, such as those resembling non-native writing patterns.
(3) PuncAttack resulted in high identification accuracy, which is expected—as even minimal edits like punctuation noise become salient when directly compared to a clean version, making extraneous or altered characters visually stand out. However, its moderate localization accuracy suggests that while readers often sense something is off, they struggle to pinpoint the exact manipulated paragraph—especially when such subtle noise is sparsely distributed across a lengthy article.
Besides, in real-world scenarios without prior knowledge of an attack, such perturbations may be easily dismissed as natural typos or stylistic quirks, allowing them to evade scrutiny. Thus, despite being simple, PuncAttack retains strong potential as a subtle and deceptive adversarial strategy.
(4) TextFooler and DeepWordBug both occupy a middle ground but differ in performance characteristics. DeepWordBug does not introduce obvious special characters like punctuation, making its subtle character-level perturbations difficult to identify during comparisons of lengthy articles, which results in relatively low identification accuracy.
However, when annotators are informed about the presence of attacks and carefully examine the text, they achieve a high localization accuracy (96%), indicating that these spelling changes can be discovered with focused reading. In contrast, TextFooler exhibits lower identification and localization accuracies, suggesting that synonym substitution attacks are harder to detect due to their higher semantic preservation, which reduces their perceptibility.
These findings highlight that the detectability of adversarial attacks varies significantly depending on their nature and granularity. Attacks that preserve semantics tend to be more stealthy, while those introducing stylistic or punctuation anomalies are easier to spot in direct comparisons but still pose risks in real-world, zero-knowledge scenarios. This underscores the ongoing challenge in designing robust defenses against diverse and evolving adversarial strategies.